A HIKER WALKS 3 MILES PER HOUR. HOW MANY HOURS DOES IT TAKE THE HIKER TO WALK n MILES?
Answer by gonzo(654)

Rate times time = distance
let R = rate in miles per hour.
let T = time in hours.
let D = distance in miles.
-----
let R = 3 miles per hour.
-----
start with D = R * T
divide both sides of the equation by R
equation becomes:
D/R = T
which is the same as:
T = D / R
-----
equation we want to work with is: T = D / R
since R = 3, equation becomes:
T = D / 3
if we let D = n, equation becomes:
T = n / 3
since T = number of hours it takes for hiker to get there, equation becomes:
number of hours it takes for hiker to get there = n / 3
-----
answer is:
number of hours it takes for hiker to walk n miles is n / 3 hours.
-----
example:
let n = 300 miles
if the hiker is traveling at 3 miles per hour, it will take 300 / 3 = 100 hours to get there
-----